OpenStudy (anonymous):

i got the fraction part right but i cant figure out how to get the variables: 4m^3/3m^2n multiplied by 9n^2/30m^2n^2 = 2/5

OpenStudy (saifoo.khan):

\[\Large \frac{x^a}{x^b} \to x^{a-b}\]

OpenStudy (saifoo.khan):

\[\Large x^a \times x^b \to x^{a +b}\]

OpenStudy (anonymous):

m=1 n=1

OpenStudy (anonymous):

I'm not sure though, because it's a weird answer, let me check over my work.. it's a bit messy.

OpenStudy (anonymous):

how would that work if the variables cant be divided by each other?

OpenStudy (anonymous):

\[(4m^{3}/3m^{2}n) \times (9n^{2}/30m^{2}n^{2}) = 2/5\]

OpenStudy (anonymous):

It makes sense though... use cross cancellation, that's the method I used: you would get 2/5mn=2/5 so basically you know that: mn=1 therefor m and n must both be 1. If you plug in one as the variables... it works out! so the answer is: m=1 n=1

OpenStudy (anonymous):

im still confused. would you cancel them out going across or up and down?

OpenStudy (anonymous):

You can cross cancel things every way except across.

OpenStudy (anonymous):

that means things that are diagonal can cross out.

OpenStudy (anonymous):

can you draw out how you did that?

OpenStudy (bahrom7893):

was the question this?